OpenStudy (anonymous):

The average of x numbers is 12. If the avg of these x numbers and 4 and 6 is 10, what is the value of x?

OpenStudy (anonymous):

x= total amount of numbers lets denote x1, x2, x3 as the actual numbers and then denote xs as the sum of x1,x2,x3 then we know \[\frac{x_1 +x_2+x_3 + ...}{x} = 12\] \[\frac{x_s}{x} = 12\] now, is we add 2 more numbers, the total amount of numbers of the average would be "x +2" and the actual numbers will have 4 and 6 added in so the average equation will then look like \[\frac{x_s +4+6}{x+2}=10\]

OpenStudy (anonymous):

at this point, you have 2 equations and 2 unknowns which you can then solve by using which ever method you'd like substitution method seems like a good choice

OpenStudy (anonymous):

thank you so much. the answer is 5

OpenStudy (anonymous):

i havent actually dont the problem myself so i'll take your word for it

OpenStudy (anonymous):

yup looks right, hopefully i didnt screw up the calculation in my head